# A comparative analysis of rank aggregation methods for the partial label ranking problem

### v3

@@ 3.1 Ranking
A ranking represents a preference relation among a set of items =$\mathcal{I}=\{1,\dots,n\}$. Following standard conventions in preference learning, the pairwise preference relations between two items $u,v\in\mathcal{I}$ are denoted as $u\succ v$ if $u$ is preferred over $v$, $u\prec v$ if $v$ is preferred over $u$, and $u\sim v$ if there is no preference between $u$ and $v$.
@@ 3.3 Partial label ranking problem
Formally, the partial label ranking problem involves a finite set of $n$ items to be ranked =$\mathcal{I}=\{1,\dots,n\}$ and a set of $N$ training pairs $\mathcal{D}=\{(x^{i},\pi^{i})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{S}_{\mathcal{I}}^{\succeq,\mathrm{inc}}$. Here, $\mathcal{X}$ denotes the instance space (typically a subset of $\mathbb{R}^{m}$) and $\mathcal{S}_{\mathcal{I}}^{\succeq,\mathrm{inc}}$ denotes the set of (possibly incomplete) partial rankings defined over $\mathcal{I}$. Each pair $(x^{i},\pi^{i})\in\mathcal{D}$, consists of a configuration of values over $m$ predictive features, i.e., $x^{i}=(x_{1}^{i},\dots,x_{m}^{i})$, and a (possibly incomplete) ranking with ties of the values in $\mathcal{I}$, i.e., $\pi^{i}$. Given these, a partial label ranking algorithm learns a preference model, which is then utilized to assign a bucket order $\mathcal{B}$ to any input instance $x\in\mathcal{X}$.
@@ 5.1 Methodology
The metric of interest is the $\tau_{X}$ rank correlation coefficient [22]. Similar to how Kendall’s $\tau$ is used for label ranking [15], $\tau_{X}$ is specifically designed to deal with tied rankings. As a result, it is a valuable and widely adopted metric for evaluating partial label ranking performance [5]. $\tau_{X}$ ranges from 1-1- 1 (complete disagreement) to 111 (perfect agreement), with 0 indicating no correlation between the predicted and true bucket orders. Algorithms were assessed using five repetitions of a ten-fold (5×105$\times$ 105 × 10) cross-validation method. To test robustness to missing information, 0%, 30%, and 60% of class labels were randomly omitted. These percentages are standard values used for label and partial label ranking problems [5, 7, 16]. Finally, the results were analyzed following the procedure in [19, 27] using the exreport tool [10]. A Friedman test [26] was conducted to determine whether the performance differences among algorithms are statistically significant. This non-parametric test ranks algorithms across datasets and checks if the rank differences are greater than expected by chance. If the null hypothesis is rejected, a Holm’s post-hoc test [30] is applied to compare the top-ranked method against the others, while controlling the family-wise error rate. The Friedman test and the post-hoc analysis were performed at a 5% significance level.
@@ 5.3 Learning methods
Figure 2 shows the impact of the $\beta$ hyperparameter on the $\tau_{X}$ score for the Borda and Copeland algorithms for varying percentages of missing class labels. Notably, the best $\beta$ value is 0.9 for Borda and 0.4 for Copeland, regardless of the percentage of missing class labels. As $\beta$ increases, both methods become more likely to group items into the same bucket, with the extreme case being a single bucket for all items. For Copeland, this critical point occurs at $\beta=0.5$ because the scoring function assigns all items a score of 0.5. This is because the entire pair order matrix falls within the interval $[0.5-\beta,0.5+\beta]=[0,1]$, matching its full value range, as discussed in Section 3.2. Therefore, in the rest of the experimental evaluation, we set $\beta=0.9$ for Borda and $\beta=0.4$ for Copeland.
@@ 5.4.2 Real-world datasets
We trained a linear regression model using the following dataset meta-features as predictors for the new datasets: number of instances, number of features, number of classes, average number of buckets, and average percentage of missing class labels. Thus, each training sample represents a unique combination of a dataset and a specific percentage of missing class labels, with percentages ranging from 0 to 60 in increments of 10. These averages were computed from the 5×105$\times$ 105 × 10 cross-validation method carried out. The target variable for the model was the $\beta$ yielding the best average results.

### v4

@@ 3.1 Ranking
A ranking represents a preference relation among a set of items $\mathcal{I}=\{1,\dots,n\}$. Following standard conventions in preference learning, the pairwise preference relations between two items $u,v\in\mathcal{I}$ are denoted as $u\succ v$ if $u$ is preferred over $v$, $u\prec v$ if $v$ is preferred over $u$, and $u\sim v$ if there is no preference between $u$ and $v$.
@@ 3.3 Partial label ranking problem
Formally, the partial label ranking problem involves a finite set of $n$ items to be ranked $\mathcal{I}=\{1,\dots,n\}$ and a set of $N$ training pairs $\mathcal{D}=\{(x^{i},\pi^{i})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{S}_{\mathcal{I}}^{\succeq,\mathrm{inc}}$. Here, $\mathcal{X}$ denotes the instance space (typically a subset of $\mathbb{R}^{m}$) and $\mathcal{S}_{\mathcal{I}}^{\succeq,\mathrm{inc}}$ denotes the set of (possibly incomplete) partial rankings defined over $\mathcal{I}$. Each pair $(x^{i},\pi^{i})\in\mathcal{D}$, consists of a configuration of values over $m$ predictive features, i.e., $x^{i}=(x_{1}^{i},\dots,x_{m}^{i})$, and a (possibly incomplete) ranking with ties of the values in $\mathcal{I}$, i.e., $\pi^{i}$. Given these, a partial label ranking algorithm learns a preference model, which is then utilized to assign a bucket order $\mathcal{B}$ to any input instance $x\in\mathcal{X}$.
@@ 5.1 Methodology
The metric of interest is the $\tau_{X}$ rank correlation coefficient [22]. Similar to how Kendall’s $\tau$ is used for label ranking [15], $\tau_{X}$ is specifically designed to deal with tied rankings. As a result, it is a valuable and widely adopted metric for evaluating partial label ranking performance [5]. $\tau_{X}$ ranges from 1-1 (complete disagreement) to $1$ (perfect agreement), with 0 indicating no correlation between the predicted and true bucket orders. Algorithms were assessed using five repetitions of a ten-fold ($5\times 10$) cross-validation method. To test robustness to missing information, 0%, 30%, and 60% of class labels were randomly omitted. These percentages are standard values used for label and partial label ranking problems [5, 7, 16]. Finally, the results were analyzed following the procedure in [19, 27] using the exreport tool [10]. A Friedman test [26] was conducted to determine whether the performance differences among algorithms are statistically significant. This non-parametric test ranks algorithms across datasets and checks if the rank differences are greater than expected by chance. If the null hypothesis is rejected, a Holm’s post-hoc test [30] is applied to compare the top-ranked method against the others, while controlling the family-wise error rate. The Friedman test and the post-hoc analysis were performed at a 5% significance level.
@@ 5.3 Learning methods
Figure 2 shows the impact of the $\beta$ hyperparameter on the $\tau_{X}$ score for the Borda and Copeland algorithms for varying percentages of missing class labels. Notably, the best $\beta$ value is 0.9 for Borda and 0.4 for Copeland, regardless of the percentage of missing class labels. As $\beta$ increases, both methods become more likely to group items into the same bucket, with the extreme case being a single bucket for all items. For Copeland, this critical point occurs at $\beta=0.5$ because the scoring function assigns all items a score of 0.5. This is because the entire pair order matrix falls within the interval $0.5-\beta,0.5+\beta]=[0,1]$, matching its full value range, as discussed in Section 3.2. Therefore, in the rest of the experimental evaluation, we set $\beta=0.9$ for Borda and $\beta=0.4$ for Copeland.
@@ 5.4.2 Real-world datasets
We trained a linear regression model using the following dataset meta-features as predictors for the new datasets: number of instances, number of features, number of classes, average number of buckets, and average percentage of missing class labels. Thus, each training sample represents a unique combination of a dataset and a specific percentage of missing class labels, with percentages ranging from 0 to 60 in increments of 10. These averages were computed from the $5\times 10$ cross-validation method carried out. The target variable for the model was the $\beta$ yielding the best average results.